OpenStudy (anonymous):

Elimintation 2x+y=104 17x-2y=464

OpenStudy (mertsj):

1. Multiply the top equation by 2 2. Add that answer to the second equation 3. Divide both sides of that equation by 21

OpenStudy (anonymous):

I got 17x-2y=464 2x+2y=208 19x=672

OpenStudy (anonymous):

I get a decimal answer and i think thats wrong

OpenStudy (anonymous):

\[2x+y=104\\ 17x-2y=464\]double the first one get \[4x+2y=208 \\ 17x-2y=464\]

OpenStudy (anonymous):

then when you add you get \[21x=672\]

OpenStudy (anonymous):

you good from there?

OpenStudy (anonymous):

X=32 God I make so many stupid small mistakes

OpenStudy (anonymous):

annoying, aren't they?

OpenStudy (anonymous):

Completely -__- can you stick aroun so I can show you my y?

OpenStudy (anonymous):

y=40